OpenStudy (anonymous):

SIMPLIFY THE EXPONENTIAL ESPRESSION...HELP:( X^10 y^0

OpenStudy (anonymous):

\(y^0=1\)

OpenStudy (anonymous):

X^10

OpenStudy (anonymous):

because like satellite posted. y^0=1

OpenStudy (anonymous):

sorry to be complicated but, how do you solve it?

OpenStudy (anonymous):

\[y^0=1\] \[x^{10} \times 1=?\]

OpenStudy (anonymous):

x^10?

OpenStudy (anonymous):

okay thankyou!

OpenStudy (anonymous):

yes. you're welcome.

OpenStudy (anonymous):

so would (-12x8y/678xy)^0 =1?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

anything to the exponent of 0 equals 1

OpenStudy (anonymous):

okay so how about..... 9x^-2y^8

OpenStudy (anonymous):

what's the exponent of 9x? is it the whole (-2y^8)?

OpenStudy (anonymous):

the exponent of 9x is -2

OpenStudy (anonymous):

it would equal \[9y^8 \over x^2\]

OpenStudy (anonymous):

because usually you don't leave negative exponents as the answer. To make it positive, you make the unknown the denominator and it becomes positive.

OpenStudy (anonymous):

okay so 9x^-2 would be simplified to what? because i know you have to make it a positive number and then multiply it with y^8

OpenStudy (anonymous):

Here's a link you can check out. http://www.purplemath.com/modules/exponent2.htm

OpenStudy (anonymous):

\[9y^8 \over x^2\] That's the answer. Read the link for a more detailed explanation.

OpenStudy (anonymous):

ok thankyou!!!!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

sorry to keep bothering you but what about (-5x^5y^-6)(2x^-1y)

OpenStudy (anonymous):

you have to multiply them.